OpenStudy (anonymous):

a car of mass 1000kg travelling along a straight horizontal road accelerates uniformly from 15m/s to 25m/s in a distance of 320m. if the resistance to motion is 145N,find the driving force of the engine.

OpenStudy (ash2326):

Lianaro it'd help you more, if you post this in physics. You'd get better response there. Thanks :)

OpenStudy (ash2326):

Nevertheless, I'm solving it for you here :)

OpenStudy (anonymous):

thnx :)... this question actually in my tutorial

OpenStudy (ash2326):

Mass=1000 Kg it's given car accelerates from 105 to 25 in a distance of 320 m we know \[v^2-u^2=2as\] v=final velocity= 25 m/s u=initial velocity= 15m/s s=320 so \[625-225=2*320*a\] a=400/640 m/s^2 or a=5/8m/s^2 Force= m*a now Force due to this =100*5/8=62.5 Newtons it's given the resistance is 145 N so the driving force is= 145+62.5=207.5 Newtons

OpenStudy (anonymous):

tq :)